two walls of a house have windows one each of the shape of a rectangle having a semicicle on the top the rectangle part is 1.2 m long and 4.2 cm wide from how much area the air can come in the room
Answer by Cromlix(4381)

Hi there,
Consider one window:
area of a rectangle = length x width
Area = 1.2 x 4.2 = 5.04 m^2
area of semicircle = (pi x radius^2)/2
Area = (pi x 2.1^2)/2 = 6.93 m^2 (2 decimal places)
Total area = 5.04 + 6.93 = 11.97 m^2
2 windows = 2 x 11.97 m^2
= 23.94 m^2
Hope this helps :-)